### v1

# MLLM-Fabric: Multimodal Large Language Model-Driven Robotic Framework for Fabric Sorting and Selection

Abstract
Choosing the right fabric is crucial to meet functional and quality requirements in robotic applications for textile manufacturing, apparel production, and smart retail. We present MLLM-Fabric, a robotic framework powered by multimodal large language models (MLLMs) for fabric sorting and selection. The system includes a robotic arm, a camera, a visuotactile sensor, and a pressure sensor. It employs supervised fine-tuning and multimodal explanation-guided knowledge distillation to accurately classify and rank fabric properties. To facilitate further research, we release a dataset of 220 unique fabric samples, including RGB images and synchronized visuotactile and pressure data. Experimental results show that our Fabric-Llama-90B model consistently outperforms pretrained vision-language baselines in both property ranking accuracy and selection reliability.
I Introduction
Robotic fabric manipulation has seen growing interest in recent years, especially for tasks involving material identification, physical property estimation, and deformable object handling [1]. While visual and tactile sensing have improved perception capabilities, conventional classification methods often fail to meet real-world selection demands. For instance, fabrics labeled under the same category, such as “cotton,” may vary drastically in softness, elasticity, or thickness. These differences critically affect performance in downstream tasks such as folding, dressing, or tool interaction. Effective fabric selection requires reasoning over continuous material attributes and contextual task relevance. Visual-only methods struggle to capture compliance or friction-related cues, while existing multimodal approaches [2, 3, 4, 5] often rely on unimodal embeddings, lack physical property supervision, or treat fabrics as static object categories without supporting structured comparison or functional decision-making. These limitations underscore the need for new paradigms in robotic material understanding.
This work introduces a physically grounded and task-aware robotic framework for fabric selection, formulated as a property-specific pairwise comparison problem. Instead of assigning absolute class labels, the system learns to determine which of two materials better satisfies a given property, such as softness or elasticity. This relative reasoning supports nuanced, context-sensitive decisions that are critical in real-world applications like adaptive grasping or clothing recommendation, where functional behavior rather than class identity determines success. At the core of our system is a multimodal language model trained on synchronized RGB, tactile, and pressure data. It is supervised through pairwise preference learning to support physically grounded, interpretable comparison of material properties. We further apply explanation-guided distillation to enhance reasoning clarity and allow abstention in uncertain cases. To support this formulation, we construct a dataset of 220 real-world fabrics, each annotated with four physical attributes and paired with co-registered RGB images, GelSight sequential frames, and pressure values. This dataset enables reproducible benchmarking of property-level material comparison under realistic sensory conditions.
Figure 1: The robotic system uses multimodal perception and reasoning to select fabrics based on user needs, choosing pure cotton seersucker (the white one on the hanger) as the optimal summer clothing. GelSight images are sequential frames with corresponding pressure sensor data. The figure shows a pressure trend graph for one fabric as an example.
Figure 2: Fabric selection robotic system workflow. The system follows a four-stage pipeline. First, a Visual Perception Module captures fabric IDs, RGB images, and 3D positions. Second, a Tactile Perception Module records GelSight image sequences and pressure values. Third, an MLLM Reasoning Module processes multimodal data to compare fabrics. Finally, a Task Execution Module selects and retrieves the goal fabric using stored IDs and positions. Fabric data is archived for efficient decision-making.
The main contributions of this work are as follows:
  • A new problem formulation that recasts robotic material selection as a property-specific pairwise comparison task, enabling functional decision-making under ambiguity and task variability.
  • A multimodal large language model framework trained using pairwise preference supervision and explanation-guided knowledge distillation, which supports interpretable and abstention-aware predictions from tactile, visual, and force inputs.
  • A robotic system that performs property-level reasoning for deformable object selection in tasks where relative material behavior, such as softness or elasticity, is more informative than discrete class identity.
  • A real-world dataset of 220 fabrics with co-registered RGB, GelSight, and pressure data, annotated with semantic physical properties, to support future research in semantic perception and robotic material understanding.
A demonstration video is included in the submission. The code and full dataset will be made publicly available on HuggingFace and a project website upon publication.
II Related Work
While prior work has explored visuotactile integration, most approaches treat fabric understanding as a classification or retrieval task. In contrast, our work frames it as property-specific, pairwise reasoning for task-conditioned selection, requiring different forms of supervision and inference. Yuan et al. [2] trained CNNs on RGB, depth, and GelSight data to learn modality-invariant embeddings for individual fabrics, enabling cross-modal inference but not supporting comparison or task-driven selection. Their subsequent work [3] developed an autonomous robot system to classify fabric properties using tactile images collected via Kinect-guided squeezing, but still lacked multimodal fusion or pairwise reasoning. ViTac [4] used deep maximum covariance analysis to embed visual and tactile data into a shared latent space for texture classification, but did not address functional or mechanical properties. Cao et al. [5] proposed a generative method to synthesize tactile features from vision and language, enabling zero-shot recognition, but only produced coarse semantic labels without supporting fine-grained attribute ranking or selection.
Recent work integrates tactile inputs into large-scale vision-language frameworks. TVL [6] trained a touch-vision-language encoder on pseudo-labeled data for open-vocabulary material recognition. However, it lacks physical supervision, material comparison, or structured selection capabilities. UniTouch [7] aligned multiple tactile sensors with vision-language spaces, supporting cross-modal generalization, but omitted property supervision and any form of reasoning or ranking. Among existing methods, Octopi [8] is the most closely related to our work. It introduces visuotactile prompting for property inference, but operates at the object level by predicting coarse attributes of entire items. In contrast, our framework performs material-level reasoning over localized fabric properties and supports task-conditioned comparison and selection.
Table I summarizes eight representative methods across six functional axes relevant to material selection. Our method is the only one to unify all six capabilities in a fabric-specialized system. These axes span from task scope (fabric specificity), to perceptual coverage (vision-touch fusion), to reasoning ability (property comparative inference and ranking), and finally to task conditioning and language-based generalization. By framing material understanding as a grounded, pairwise reasoning task enriched with explanation generation, our system bridges the gap between semantic perception and task-driven robotic manipulation, enabling interpretable and goal-aligned material decisions.
TABLE I: Comparison of representative methods across capabilities: fabric-specific focus (Fabric), multimodal fusion (MM), comparative reasoning (CR), ranking (RK), task awareness (TA), and use of multimodal language models (MLLM). Refers to the study titled Active Clothing
Method Fabric MM CR RK TA MLLM
Yuan et al. [2]
Yuan et al. [3]
ViTac [4]
Cao et al. [5]
TVL [6]
UniTouch [7]
Octopi [8]
Ours
III Methodology
This section describes the methodology of the proposed fabric selection robotic system (Fig. 1), detailing the data acquisition, multimodal perception, model training, and inference processes that enable accurate fabric property prediction and selection.
III-A Visual Data Processing and Fabric Localization
The overall system workflow is illustrated in Fig. 2. During system operation, an Intel RealSense D435 camera mounted on the robotic arm captures RGB images and depth data to provide the necessary visual information for fabric localization and segmentation. To facilitate fabric identification and ensure consistent data association, ArUco tags are attached to each fabric sample, serving as unique IDs for pose tracking on the hanger. This step uses standard computer vision techniques to identify and record each sample’s position and orientation in 3D space. Depth data along the Z-axis is integrated with ArUco-based tracking to enhance pose estimation and account for shape variations or hanger misalignments. This integrated process ensures reliable visual localization, supporting the subsequent multimodal perception pipeline.
Following localization, the Segment Anything Model (SAM) [9] performs pixel-level segmentation to isolate each fabric sample from the background and neighboring objects in the RGB image. The segmented images are resized to 224$\times$224 pixels and stored with structured metadata, comprising tuples $(F_{i},P_{i},I_{i})$, where $F_{i}$ denotes the fabric ID, $P_{i}\in\mathbb{R}^{3}$ represents the 3D position vector, and $I_{i}$ corresponds to the segmented RGB image. This structured data ensures consistency across sensing modalities and supports downstream multimodal reasoning and property prediction tasks.
Figure 3: (A) Overview of the tactile sensing setup, including the GelSight Mini sensor and a planar pressure sensor mounted on a hard surface. (B) The layered structure during pressing: (a) gel layer, (b) fabric sample, (c) hard surface, and (d) pressure sensor. (C) Synchronization of tactile image sequences from the GelSight sensor with corresponding pressure measurements at a frequency of 25 Hz. (D) A selection of ten representative fabric samples from a dataset containing 220 unique fabrics.
III-B Fabric Tactile Sensing and Force-labeled Tactile Data
A dual-sensor setup is used for fabric tactile sensing (Fig. 3A): the left fingertip of the Franka Hand mounts a GelSight Mini [10] for high-resolution visuotactile imaging, while the right fingertip holds a strain-based planar pressure sensor with a dome surface for direct force measurement. The layered structure appears in Fig. 3B.
Although the GelSight sensor can estimate force from surface deformation, its accuracy degrades significantly on fabrics due to sensitivity to local thickness, compliance [3], and device-level variation [10]. To avoid relying on such uncertain estimates, we use a strain-gauge pressure sensor (HZC-30B, Chengying Sensor, China) mounted beneath a rigid planar surface. The sensor has a rated range of 0–10 kg, sensitivity of 1.0–1.5 mV/V, nonlinearity within $\pm$0.3full scale (F.S.), and creep within $\pm$0.5F.S. It operates from $-30\,^{\circ}\mathrm{C}$ to $70\,^{\circ}\mathrm{C}$ and supports 200% F.S. overload.
The GelSight and pressure sensor are mounted on opposing gripper fingertips to compress the fabric simultaneously. By Newton’s third law, the pressure reading equals the normal contact force on the GelSight surface. This setup enables synchronized recording of force and deformation without relying on image-based inference. The resulting stream comprises force-labeled tactile data, with each GelSight frame aligned to a corresponding pressure value. We use “force” to refer to these pressure readings throughout. This alignment substitutes for GelSight force calibration and provides accurate ground-truth supervision during training.
III-C Multimodal Data Collection and Fusion Strategy
We collect data from 220 fabrics using a Franka arm, gripper, RealSense D435, GelSight Mini, and a planar pressure sensor. Each fabric is pressed twice (10s each) with a moderate-speed approach and slow closure for stable contact. GelSight and pressure data are sampled at 25 Hz and time-synchronized (Fig.3C), pairing each frame with a force reading. Fig.3D shows RGB examples illustrating visual diversity. The dataset contains RGB images, GelSight sequences, and aligned pressure values. GelSight captures spatial deformation linked to texture and compliance but not force; the pressure sensor provides scalar readings. Fusion preserves both deformation context and contact intensity.
To integrate this heterogeneous information, we first extract feature embeddings from each modality: RGB and GelSight images are processed by a frozen visual encoder, while pressure signals are represented in a language-compatible form and embedded via the model’s input projection layers. The resulting embeddings are unified within a shared token space and jointly processed by a multimodal transformer. This early-fusion strategy allows the model to learn cross-modal cues during training and inference, consistent with practices in recent vision-language models that align diverse modalities through token-level reasoning [11].
Each fabric sample is also annotated by experts with categorical labels for elasticity, softness, thickness, and texture. These annotations support supervised fine-tuning and evaluation.
III-D Supervised Fine-Tuning
In our study, supervised fine-tuning of an MLLM using directly labelled training data is referred to as Directly Supervised Fine-Tuning (D-SFT). Unlike knowledge distillation, which refines predictions based on teacher model outputs, D-SFT directly learns from human-annotated comparisons, ensuring explicit alignment of model parameters with expert judgments.
III-D1 Supervised Fine-Tuning Objective
The goal of supervised fine-tuning is to train the model to rank fabric properties based on pairwise comparisons. Since each fabric sample is unique, no two fabrics share identical attributes. The model’s primary task is to establish a ranking, with cases of minimal or indistinguishable differences classified as inconclusive.
The training objective is formulated as a classification task, where the model outputs a probability distribution over the three possible outcomes. The learning process is guided by a cross-entropy loss function:
$L_{\text{train}}=\sum_{i=1}^{N}\sum_{q\in Q^{\prime}}L_{\text{CE}}\left(g(z_{i},\phi),y_{i,q}\right)$ (1)
where $g(z_{i},\phi)$ denotes the predicted probability distribution for fabric property $q$, and $y_{i,q}$ represents the corresponding ground truth label. The input $z$ includes multimodal information, incorporating integrating visual, tactile, and pressure data to capture relevant fabric characteristics.
III-D2 LoRA-Based Fine-Tuning
To enable efficient and task-specific adaptation, we adopt Low-Rank Adaptation (LoRA) [12] for fine-tuning. LoRA introduces trainable low-rank matrices into the projection layers of the pre-trained language model, allowing effective parameter updates with minimal overhead. Formally, the adapted weights are defined as:
$W^{\prime}=W+\Delta W,\quad\Delta W=AB,$ (2)
where $A\in\mathbb{R}^{d\times r}$ and $B\in\mathbb{R}^{r\times d}$, with rank $r\ll d$. We explore two strategies for integrating LoRA into our framework: (1) applying fine-tuning to both the vision encoder and language module, and (2) applying LoRA-based fine-tuning solely to the language module while keeping the visual backbone frozen. The latter is motivated by findings in recent multimodal literature suggesting that pretrained visual encoders can retain strong generalization even when frozen, particularly in tasks involving limited domain-specific data [11].
III-D3 Inference Process
During inference, the model processes a given pair of fabric samples and predicts the most probable ranking for the specified property. The decision is obtained by selecting the class with the highest predicted probability:
$\hat{y}_{q}=\arg\max g(z,\phi)$ (3)
This pairwise prediction mechanism enables the model to differentiate fabric properties effectively based on direct comparisons. D-SFT directly optimizes for pairwise classification using labelled data. While knowledge distillation benefits from leveraging additional synthetic training signals, D-SFT is constrained by the quality and availability of labelled comparisons.
Figure 4: Multimodal Reasoning and Knowledge Distillation Pipeline. The upper section illustrates Multimodal Explanation-Guided Knowledge Distillation, where a teacher model generates post hoc explanations that are subsequently distilled into a student model. The lower section depicts the fabric property comparison and selection process, where the fine-tuned model performs ad hoc reasoning to compare attributes and issue fabric selection commands.
III-E Explanation-Guided Knowledge Distillation
To enhance ranking accuracy and interpretability, we introduce a Multimodal Explanation-Guided Knowledge Distillation (MEKD) process that integrates expert-validated reasoning. GPT-4o [13], being a closed-source model and significantly larger than Llama3.2-Vision-90B [14], is impractical for direct deployment due to its restricted accessibility and high computational costs. Instead, we distill GPT-4o’s reasoning ability into the more efficient Llama3.2-Vision-90B, transferring its structured justifications through MEKD while preserving interpretability. As shown in Fig. 4, this process involves pairwise fabric comparisons and AI-generated rationales from GPT-4o, which are directly incorporated into the training pipeline to enhance model supervision and improve the reliability of the distilled model.
III-E1 Post Hoc Reasoning
The teacher model first generates post hoc ranking explanation content based on ground truth data. Human experts refine these explanations to improve clarity and reliability before using them to fine-tune the student model. This process is formulated as:
$Q_{\text{post hoc}}=\text{MLLM}({z},Q_{\text{inst}},\tilde{r})$ (4)
where MLLM($z,Q^{\text{inst}},r$) denotes the Multimodal Large Language Model module that generates post hoc ranking explanations based on the input pair $z$, the instruction $Q^{\text{inst}}$, and the expert-validated ranking $r$. This approach enhances interpretability, aligns model predictions with expert knowledge, and provides structured supervision to improve generalization.
III-E2 Ad Hoc Reasoning
In the absence of ground truth labels, the student model generates ad hoc reasoning to approximate expert explanations. To enforce consistency, we minimize the discrepancy between post hoc and ad hoc outputs via a cross-entropy loss:
$\hat{\phi}=\arg\min_{\phi}\mathcal{L}_{\text{CE}}(Q_{\text{ad hoc}},Q_{\text{post hoc}})$ (5)
where $Q_{\text{ad hoc}}$ represents the model’s self-generated reasoning content. By aligning self-generated justifications with expert-reviewed rationales, the model improves reasoning consistency and robustness, enabling adaptation to unseen data.
IV Experimental Setup
This section describes the experimental configuration employed to assess the effectiveness of the proposed methods.
IV-A Fabric Selection Scenario Design and Evaluation
We evaluate the models’ ability to reason about fabric properties in ten expert-designed selection scenarios, covering diverse applications such as apparel, home textiles, and industrial fabrics. Each scenario is designed to assess the model’s ability to differentiate fabrics based on attributes like elasticity, softness, thickness, and texture, requiring both simple and multi-attribute reasoning.
We compare Fabric-Llama (Llama-Vision-90B fine-tuned with MEKD) against the original Llama-Vision-90B and GPT-4o. In these unseen scenario evaluations, models analyze multimodal data to recommend the most suitable fabric based on expert-defined criteria. Accuracy is measured by comparing their selections with expert judgments. The scenarios are structured to reflect real-world selection challenges, with increasing complexity when properties are closely ranked or when multiple attributes must be considered simultaneously. This setup evaluates whether fine-tuning enhances the model’s ability to generalize and make precise selections beyond basic classification.
IV-B Model Training and Evaluation Setup
Each fabric sample is annotated with a four-dimensional attribute vector $\mathbf{a}=[a_{1},a_{2},a_{3},a_{4}]\in\{0,1,2\}^{4}$, representing elasticity, softness, thickness, and surface texture. Attributes are discretized into three ordinal levels: 0 (low), 1 (moderate), and 2 (high). To enhance supervision clarity, we retain only samples labeled 0 or 2 for training, discarding moderate-level examples to avoid semantic ambiguity and improve contrast in pairwise learning.
Two evaluation settings are used. The seen-fabric setting selects 200 training fabrics to form 24,000 extreme-level comparison pairs. From these, 4,000 pairs are randomly sampled for training and 400 for validation, ensuring equal distribution across the four attributes. Fabric instances may recur, but pairwise combinations are not duplicated across sets. The unseen-fabric setting evaluates generalization using 20 additional fabrics entirely excluded from training. Pair construction follows the same protocol with strict fabric-level isolation to prevent overlap. To assess extrapolation beyond binary extremes, we further evaluate the model on fine-grained comparisons involving adjacent attribute levels, such as 0 versus 1 and 1 versus 2. These cases are excluded from training and used solely to test the model’s ability to reason over subtle differences.
To mitigate overcommitment in ambiguous comparisons [15], we incorporate an abstain option into all prompts. While this mechanism is designed to improve reliability under uncertainty, its impact is not the focus of our primary evaluation. As an initial exploration, we conduct a separate test on 50 curated duplicate fabric pairs excluded from training. Our fine-tuned models abstained in fewer than 50 percent of these cases, whereas GPT-4o abstained in over 90 percent, indicating more conservative decision behavior. These findings highlight the limitations of our current training setup and motivate future work on explicit duplicate-aware supervision.
We evaluate Llama3.2-Vision-90B, chosen for its strong multimodal reasoning, and DeepSeek Janus-Pro-7B, a lighter baseline. GPT-4o is excluded due to its proprietary nature and reliance on closed APIs, which hinder reproducibility and integration into real-time systems. Fine-tuning is conducted on high-performance hardware. Llama3.2-Vision-90B is trained on an NVIDIA A800 GPU (80 GB RAM) for 6.5 hours. Janus-Pro-7B and Llama3.2-Vision-11B are trained on an NVIDIA A6000 (48 GB RAM) for 3 and 5 hours, respectively.
IV-C Evaluation Metrics
We use three metrics to evaluate model performance: Property-wise Accuracy, Average Accuracy, and Prediction Skewness. Property-wise Accuracy (ACC) measures correctness for each fabric property and is defined as:
$\text{ACC}=\frac{TP+TN}{TP+TN+FP+FN},$ (6)
where $TP$, $TN$, $FP$, and $FN$ denote true/false positives and negatives, respectively. Mean Accuracy (mean ACC) is the mean ACC across the four fabric properties and reflects overall ranking performance. To quantify prediction balance, we define Prediction Skewness (SK) as:
$\text{SK}=2\times\left|p_{0}-0.5\right|\times 100,$ (7)
where $p_{0}$ is the proportion of predictions assigned to category 0. A SK of 0 indicates perfectly balanced outputs; higher values indicate increasing bias toward one class. Since ground-truth labels are evenly distributed (50% per class), lower SK is preferred as it reflects greater prediction balance.
V Experimental Results
Figure 5: Ablation study of input modalities for pairwise property comparison using GPT-4o (top) and Llama3.2-Vision-90B (bottom). Bars show attribute-wise accuracy (%), dashed lines indicate mean accuracy, and dotted lines show prediction skewness (SK; lower is better). Error bars denote standard error over three runs. “Single GelSight” = single-frame GelSight; “GelSight Seq” = Gelsight sequential frames.
Model Fine-tuning Elasticity ACC $\uparrow$ Thickness ACC $\uparrow$ Texture ACC $\uparrow$ Softness ACC $\uparrow$ mean ACC $\uparrow$ SK $\downarrow$
DeepSeek Janus-Pro-7B None 28 46 40 41 39 84.4
DeepSeek Janus-Pro-7B D-SFT 50 50 50 50 50 100
Llama3.2-Vision-11B None 53 46 44 53 49 60
Llama3.2-Vision-11B D-SFT 50 51 45 42 47 31.4
Llama3.2-Vision-90B None 46 56 52 35 47 2.1
Llama3.2-Vision-90B D-SFT 95 100 95 82 93 1
Fabric-Janus-7B MEKD 50 50 50 47 47 5.5
Fabric-Llama-11B MEKD 51 59 52 50 53 33.1
Fabric-Llama-90B* MEKD (full fine-tuning) 55 93 81 62 73 4.5
Fabric-Llama-90B MEKD (frozen vision encoder) 90 100 97 89 94 0
TABLE II: Attribute-wise evaluation of different models trained on RGB, GelSight sequential frames, and synchronized force input for pairwise fabric property comparison. The table reports agreement accuracy with expert judgments on elasticity, thickness, texture, and softness, as well as mean accuracy and prediction skewness (SK). Asterisk (*) indicates full fine-tuning of both vision and language modules; models without * use MEKD with the vision encoder frozen.
Model Original Test Fine-grained Test Unseen Fabrics
mean ACC $\uparrow$ SK $\downarrow$ mean ACC $\uparrow$ SK $\downarrow$ mean ACC $\uparrow$ SK $\downarrow$
Llama3.2-Vision-90B 47 5.5 39.7 12.8 42.8 13.6
Llama3.2-Vision-90B + D-SFT 93 1 51.7 55.5 55.7 7.4
Fabric-Llama-90B 94 0 52.2 54.5 65.7 1.8
TABLE III: Generalization performance of selected models on three evaluation settings: (1) original test pairs from seen fabrics, (2) fine-grained comparisons with subtle attribute differences (level 1 vs. 0 or 2), and (3) comparisons on unseen fabrics. Metrics include agreement accuracy with expert pairwise rankings and prediction skewness (SK). All models use RGB, GelSight sequential frames, and force-labeled inputs.
We evaluate MLLM-Fabric across modalities, learning strategies, and model configurations, and demonstrate its advantages in property prediction, generalization, and selection tasks compared to prior baselines.
V-A Fabric Property Sorting Experiment
We analyze how data modalities, training strategies, and model scales affect the accuracy, confidence, and generalization of pairwise fabric property prediction.
V-A1 Impact of Data Modalities
We conduct a systematic ablation across three input modalities: RGB, GelSight (single-frame or 4-frame sequences), and synchronized force labels. All meaningful combinations are evaluated to quantify each modality’s contribution to pairwise fabric property comparison. As shown in Fig. 5, the best performance is achieved using the full combination of RGB, sequential GelSight, and force input, reaching 65.2% accuracy for GPT-4o and 47.6% for Llama3.2. The poorest results occur with unlabeled GelSight sequences, yielding 19.4% for GPT-4o and 30.8% for Llama3.2. These findings indicate that sequential GelSight frames capture richer tactile cues than single-frame inputs, and that explicit force supervision is more critical for reliable multimodal reasoning than merely increasing input diversity.
V-A2 Impact of Training Methodologies
We compare two training paradigms: Directly Supervised Fine-Tuning (D-SFT), which optimizes model predictions based on expert-labeled pairwise comparisons, and Multimodal Explanation-Guided Knowledge Distillation (MEKD), which transfers structured rationales from a larger teacher model. As shown in Table II, D-SFT substantially improves the performance of Llama3.2-Vision-90B (mean ACC: 93%, SK: 1), and MEKD further enhances both accuracy (94%) and prediction balance (SK: 0), indicating stronger generalization and reduced bias. We also evaluate full fine-tuning of both the vision encoder and the language module. This configuration achieves a mean ACC of 73% and a skewness of 4.5, indicating lower accuracy and poorer output balance compared to the frozen-encoder variant. While less effective under limited data, full fine-tuning may still be useful given larger datasets or stronger regularization. For the smaller Janus-Pro-7B, D-SFT results in highly imbalanced outputs (SK: 100), while MEKD reduces skewness to 5.5, confirming its effectiveness in improving prediction stability even for compact architectures.
V-A3 Impact of Model Scale
Llama3.2-Vision-90B consistently outperforms Llama3.2-Vision-11B, reaching a mean ACC of 93–94% compared to sub-60% for the smaller model. For Elasticity, D-SFT yields 95% mean ACC, indicating that larger models can learn robust patterns even without post hoc reasoning. This suggests diminishing returns for post hoc supervision as model scale increases.
V-A4 Model Generalization
As discussed in Section IV.B, the model is trained using coarse-grained attribute comparisons, such as level 2 versus level 0, in order to support generalization to fine-grained distinctions, including level 2 versus level 1 and level 1 versus level 0. However, as shown in Table III, both the post hoc reasoning model and the D-SFT model perform unreliably on fine-grained comparisons, with mean ACC remaining near 50% and SK index exceeding 50, indicating high output uncertainty.
We further validate our model on a 20-sample unseen fabric split (completely excluded from training), demonstrating robust generalization by evaluating performance on 140 pair-wise comparison questions, each contrasting level-2 and level-0 samples. In this setting, the Llama3.2 baseline performs at chance level (42.8%), while the D-SFT model reaches 55.7% mean ACC with high uncertainty (SK = 7.4). Our full model (Fabric-Llama-90B) improves to 65.7% mean ACC with a reduced SK index of 1.8, indicating a moderate but meaningful generalization gain. These results suggest that while the model can extract coarse-level patterns from seen data, its generalization to novel fabrics and fine-grained distinctions remains limited. Future work should explore targeted learning strategies to handle ambiguous attribute levels and shifts in fabric identity distribution.
V-B Fabric Selection Experiment
This experiment evaluates each model’s ability to select suitable fabrics based on learned property reasoning. Unlike isolated attribute ranking, this task tests real-world decision-making. Fabric-Llama-90B, fine-tuned for this task, achieved the highest agreement with expert choices, correctly selecting 8 out of 10 fabrics (Table IV), while unfine-tuned models struggled. Selection became harder with more attributes: both failures involved three-attribute comparisons requiring finer reasoning, whereas two-attribute cases were easier. These results suggest that multi-attribute fabric selection requires structured reasoning over ranked property comparisons, with difficulty increasing as the decision involves more dimensions. Fabric-Llama-90B exhibited stepwise reasoning similar to Chain-of-Thought [16]. On an A800 GPU, sorting four fabrics took 2–3 minutes; future work will reduce inference time via parallelization and token optimization.
Scenario Properties Fabric-Llama-90B (proposed) GPT-4o Llama-Vision-90B GPT-4o IT Llama-Vision-90B IT
Sorting Acc Selection Sorting Acc Selection Sorting Acc Selection Sorting Acc Selection Sorting Acc Selection
Athletic wear
for lexibility
Elasticity (2)
Softness (2)
100% 33.30% 50% 50% 50%
Summer dresses
for children
Softness (2)
Elasticity (2)
83.30% 20% 0% 62.50% 62.50%
Light draping
fabric for
home decor
Thickness (1)
Softness (2)
66.70% 20% 50% 37.50% 37.50%
Durable fabric
for frequent-use
furniture
Texture (2)
Thickness (1)
62.50% 28.50% 62.50% 62.50% 62.50%
Protective fabric
for work
environments
Thickness (1)
Elasticity (0)
50% 50% 50% 25% 25%
Heavy-duty material
for robust
applications
Texture (2)
Thickness (2)
75% 33.30% 25% 37.50% 37.50%
Soft lining
fabric for
formal suits
Thickness (0)
Softness (2)
Texture (0)
80% 42.80% 30% 50% 50%
Lightweight smooth
fabric for
formal shirts
Thickness (0)
Softness (2)
Texture (0)
90% 45% 9.10% 75% 75%
Ultra-soft
lightweight fabric
for accessories
Softness (2)
Thickness (0)
50% 50% 16.70% 75% 75%
Thick, luxurious
fabric for
premium interiors
Thickness (2)
Elasticity (0)
Texture (2)
61.50% 23% 15.30% 41.60% 41.60%
TABLE IV: Fabric selection performance in ten real-world scenarios. The first three columns show MLLMs (Fabric-Llama-90B, GPT-4o, Llama-Vision-90B) operating on raw multimodal inputs. The last two columns represent interaction trajectory (IT) pipelines, where a CNN-LSTM baseline predicts symbolic attribute levels forwarded to a frozen MLLM. Attribute levels are encoded as 0 (low), 1 (moderate), and 2 (high). marks correct selections; indicates failures.
V-C CNN-LSTM Tactile Baseline with Interaction Trajectories
We implemented a tactile-only CNN-LSTM model [3] to classify fabrics into three levels (low, moderate, high) across four attributes. On a held-out set of 20 unseen fabrics, it achieved a mean accuracy of 58%, with lower performance on texture and softness. To evaluate symbolic reasoning, we constructed an interaction trajectory (IT) pipeline by pairing the CNN-LSTM with a non-finetuned MLLM (GPT-4o or Llama3.2-Vision-90B). Each model received only the predicted attribute levels and issued a selection decision for each scenario. As shown in Table IV, both IT pipelines solved 5 out of 10 scenarios (50% accuracy), and their predictions were highly consistent. This consistency suggests that selection outcomes were largely determined by the upstream classifier. The modular structure restricts access to raw sensory input and prevents joint reasoning over perception and decision stages. In contrast, our end-to-end MLLM framework achieved 80% accuracy, highlighting the benefit of integrated training with multimodal grounding.
These comparisons suggest that end-to-end multimodal learning is more effective than modular pipelines, especially when alignment between perception and reasoning is critical. They also highlight the limitations of using symbolic attribute prediction as a proxy, and motivate tightly coupled training for grounded selection tasks.
VI Conclusion
We present MLLM-Fabric, a robotic framework that integrates visual, tactile, and force sensing with large language model reasoning to support fabric sorting and selection. By recasting material understanding as a pairwise property comparison task, the system achieves accurate and interpretable predictions across elasticity, softness, thickness, and texture. Evaluated on 220 fabrics and ten real-world scenarios, the proposed method outperforms symbolic and pretrained baselines in both property ranking and task-specific fabric selection. This work establishes a foundation for multimodal, reasoning-capable robotic systems, with future directions including fine-grained generalization, duplicate-aware learning, and real-time decision-making for deformable object manipulation.